OpenStudy (anonymous):

Express as a logarithm of a single number or expression. 1/2(logbx - logby) I went as far as (logb x/y)^1/2 but I'm not sure if this is correct. If it is, where would I go from here?

zepdrix (zepdrix):

Be careful with your brackets. The 1/2 power is being applied to the `contents` of the log, not the log itself. \[\Large\bf\sf \frac{1}{2}\left(\log_bx - \log_by\right)\quad=\quad \frac{1}{2}\log_b\frac{x}{y}\] \[\Large\bf\sf =\quad \log_b\left(\frac{x}{y}\right)^{1/2}\]And recall that the 1/2 power is the square root, so `if you wanted to`, you could write it as,\[\Large\bf\sf \log_b\sqrt{\frac{x}{y}}\]

zepdrix (zepdrix):

Looks like you're on the right track! :) Just be careful with your brackets. That's as far as it can be simplified.

OpenStudy (anonymous):

Thank you very much! I'll make sure to watch my brackets next time.

zepdrix (zepdrix):

Oh and, \(\Large\bf \color{#008353}{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (anonymous):

Haha, thanks! :D